# The Pivoting Framework: Frank-Wolfe Algorithms with Active Set Size Control

### v2

@@ B FW variants
4       ${\mathbf{a}}^{(t)}\leftarrow\operatorname{argmax}_{{\mathbf{v}}\in{\mathcal{S}}^{(t)}}\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{v}}\rangle$triangleright away vertex
5       ${\mathbf{v}}^{(t)}\leftarrow\operatorname{argmin}_{{\mathbf{v}}\in{\mathcal{V}}}\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{v}}\rangle$triangleright FW vertex
@@
7             ${\mathbf{d}}^{(t)}\leftarrow{\mathbf{x}}^{(t)}-{\mathbf{a}}^{(t)}$triangleright away step
@@
9             $\boldsymbol{\alpha}^{(t+1)}\in\Delta_{|{\mathcal{V}}|}$ s.t. for all ${\mathbf{s}}\in{\mathcal{V}}$, $\alpha_{\mathbf{s}}^{(t+1)}\leftarrow\begin{cases}(1+\eta^{(t)})\alpha_{\mathbf{s}}^{(t)},&\text{if}\ {\mathbf{s}}\in{\mathcal{V}}\setminus\{{\mathbf{a}}^{(t)}\}\\ (1+\eta^{(t)})\alpha_{\mathbf{s}}^{(t)}-\eta^{(t)},&\text{if}\ {\mathbf{s}}={\mathbf{a}}^{(t)}\\ \end{cases}$
@@
12             ${\mathbf{d}}^{(t)}\leftarrow{\mathbf{v}}^{(t)}-{\mathbf{x}}^{(t)}$triangleright FW step
@@
14             $\boldsymbol{\alpha}^{(t+1)}\in\Delta_{|{\mathcal{V}}|}$ s.t. for all ${\mathbf{s}}\in{\mathcal{V}}$, $\alpha_{\mathbf{s}}^{(t+1)}\leftarrow\begin{cases}(1-\eta^{(t)})\alpha_{\mathbf{s}}^{(t)},&\text{if}\ {\mathbf{s}}\in{\mathcal{V}}\setminus\{{\mathbf{v}}^{(t)}\}\\ (1-\eta^{(t)})\alpha_{\mathbf{s}}^{(t)}+\eta^{(t)},&\text{if}\ {\mathbf{s}}={\mathbf{v}}^{(t)}\\ \end{cases}$
@@ 17       ${\mathbf{x}}^{(t+1)}\leftarrow{\mathbf{x}}^{(t)}+\eta^{(t)}{\mathbf{d}}^{(t)}$
4       ${\mathbf{a}}^{(t)}\leftarrow\operatorname{argmax}_{{\mathbf{v}}\in{\mathcal{S}}^{(t)}}\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{v}}\rangle$triangleright away vertex
5       ${\mathbf{w}}^{(t)}\leftarrow\operatorname{argmin}_{{\mathbf{v}}\in{\mathcal{S}}^{(t)}}\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{v}}\rangle$triangleright local FW vertex
6       ${\mathbf{v}}^{(t)}\leftarrow\operatorname{argmin}_{{\mathbf{v}}\in{\mathcal{V}}}\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{v}}\rangle$triangleright FW vertex
@@
8             ${\mathbf{d}}^{(t)}\leftarrow{\mathbf{w}}^{(t)}-{\mathbf{a}}^{(t)}$triangleright local pairwise step
@@
13             ${\mathbf{d}}^{(t)}\leftarrow{\mathbf{v}}^{(t)}-{\mathbf{x}}^{(t)}$triangleright FW step
@@
15             $\boldsymbol{\alpha}^{(t+1)}\in\Delta_{|{\mathcal{V}}|}$ s.t. for all ${\mathbf{s}}\in{\mathcal{V}}$, $\alpha_{\mathbf{s}}^{(t+1)}\leftarrow\begin{cases}(1-\eta^{(t)})\alpha_{\mathbf{s}}^{(t)},&\text{if}\ {\mathbf{s}}\in{\mathcal{V}}\setminus\{{\mathbf{v}}^{(t)}\}\\ (1-\eta^{(t)})\alpha_{\mathbf{s}}^{(t)}+\eta^{(t)},&\text{if}\ {\mathbf{s}}={\mathbf{v}}^{(t)}\\ \end{cases}$
@@ 18       ${\mathbf{x}}^{(t+1)}\leftarrow{\mathbf{x}}^{(t)}+\eta^{(t)}{\mathbf{d}}^{(t)}$
C Active set identification with the blended pairwise Frank-Wolfe algorithm
@@
$I({\mathbf{x}})=\{i\in[n]\mid\lambda_{i}({\mathbf{x}})=0\}\qquad\text{and}\qquad I^{c}({\mathbf{x}})=[n]\,\backslash\,I({\mathbf{x}}).$ (C.1)
@@
$\delta^{(t)}=\max_{i\in[n]\,\backslash\,O^{(t)}}\lambda_{i}({\mathbf{x}}^{*}),$
$\delta_{\min}=\min_{i\in I^{c}({\mathbf{x}}^{*})}\lambda_{i}({\mathbf{x}}^{*}),$ (C.2)
@@
Lemma C.1 (BPFW step multipliers).
@@
Lemma C.2 (Bounds on the multipliers).
@@
$\lambda_{i}({\mathbf{x}}^{(t)})$ $<h^{*}\left(L+\frac{\delta^{(t)}}{2}\right)$ $\mathrm{for\ all}\ i\in I({\mathbf{x}}^{*})$ (C.3)
$\lambda_{j}({\mathbf{x}}^{(t)})$ $>-h^{*}\left(L+\frac{\delta^{(t)}}{2}\right)+\delta_{\min}$ $\mathrm{for\ all}\ j\in I^{c}({\mathbf{x}}^{*}).$ (C.4)
@@
Bomze et al., (2020, Lemma 3.1) proves that $|\lambda_{i}({\mathbf{x}}^{*})-\lambda_{i}({\mathbf{x}}^{(t)})|<h^{*}\left(L+\frac{\delta^{(t)}}{2}\right)$ for any $i\in[n]$ and iteration $t$ where the assumptions of the Lemma hold. The two subsequent inequalities follow from the definitions of $I({\mathbf{x}}^{*})$ in (C.1) and $\delta_{\min}$ in (C.2). In particular, Equation (C.4) can be obtained as follows:
@@
$\eta^{(t)}\geq\min\left\{\alpha_{{\mathbf{a}}^{(t)}}^{(t)},c\frac{\langle\nabla f({\mathbf{x}}^{(t)}),-{\mathbf{d}}^{(t)}\rangle}{L\|{\mathbf{d}}^{(t)}\|_{2}^{2}}\right\}$ (C.5)
for a given $c>0$ is a sufficient condition for Assumption 1 to hold (with a modified Lipschitz constant $L/c$). This means in particular that the commonly-used adaptive step-size strategy introduced in Pedregosa et al., (2018) respects the weaker condition (C.5), with the factor $c$ corresponding to the parameter $\tau$ of the backtracking subroutine Pedregosa et al., (2018, Algorithm 2).
We can now state Theorem C.3, which proves the active set identification for BPFW. The proof follows the structure of that of Bomze et al., (2020, Theorem 3.3), where the same property is shown for AFW.
Theorem C.3 (BPFW identification).
@@
We first consider the case in which $J^{(t)}=\emptyset$. Then, $[n]\,\backslash\,O^{(t)}=I({\mathbf{x}}^{*})$ and $\delta^{(t)}=0$ since $\lambda_{i}({\mathbf{x}}^{*})=0$ for all $i\in I({\mathbf{x}}^{*})$. Inequality (C.4) becomes:
@@
Since $J^{(t)}=\emptyset$, it holds that $I^{c}({\mathbf{x}}^{*})=O^{(t)}$ and $x^{(t)}_{j}=0$ for all $j\in I^{c}({\mathbf{x}}^{*})$. Following Lemma C.1, Property 2, we conclude that $x_{j}^{(t+1)}=0$ for all $j\in I^{c}({\mathbf{x}}^{*})$. Thus, $J^{(t+1)}=\emptyset$.
We now consider the case when $|J^{(t)}|>0$ and first prove that the FW vertex $i$ and the away vertex $j$ are such that $i\in I({\mathbf{x}}^{*})$ and $j\in J^{(t)}$. Recall that ${\mathcal{S}}^{(t)}_{I}\subseteq[n]\,\backslash\,O^{(t)}=J^{(t)}\cup I({\mathbf{x}}^{*})$ and that both the FW and away vertex are selected from ${\mathcal{S}}^{(t)}_{I}$ in a pairwise step. Bomze et al., (2020, Theorem 3.3, Case 2) proves that if $|J^{(t)}|>0$, then the away vertex is always selected in $J^{(t)}$ if
@@
We next demonstrate that the FW vertex is always selected in $I({\mathbf{x}}^{*})$ and not in $J^{(t)}$. Based on (C.3) and (C.4), a sufficient condition for the FW vertex to be selected in $I({\mathbf{x}}^{*})$ is:
@@
$\geq\frac{\delta_{\min}-h^{*}(2L+\delta^{(t)})}{2L}$triangleright by subtracting (C.3) from (C.4)
@@
Theorem C.3 can be used to generalize Bomze et al., (2020, Theorem 4.3) to BPFW under similar assumptions, meaning that BPFW identifies the active set in a finite number of iterations, since the proof of Bomze et al., (2020, Theorem 4.3) does not rely on the sequence generated by the algorithm. We also note that our proof does not extend to the pairwise Frank-Wolfe algorithm (Lacoste-Julien and Jaggi,, 2015), since we rely on the BPFW property that at any iteration in which a pairwise step is performed, both the away and the FW vertex are selected from the active set.
Theorem C.4 (Active set identification property of BPFW).
@@
By affine invariance of BPFW with line-search, this is equivalent to optimizing $f(A{\mathbf{y}})$, with ${\mathbf{y}}\in\Delta_{n}$, as developed in Bomze et al., (2020, Appendix C). The proof over the simplex follows that of Bomze et al., (2020, Theorem 4.3), relying on Theorem C.3 to ensure the cardinality of the set $J^{(t)}$ decreases at each iteration once there exists a stationary point ${\mathbf{x}}^{*}$ of the problem with ${\mathbf{x}}^{*}\in{\mathcal{X}}$, such that $\|{\mathbf{x}}^{(t)}-{\mathbf{x}}^{*}\|_{1}<h^{*}$. ∎
D Algorithmic aspects
@@
E Numerical experiments (extended)
We assess our algorithm on efficiency in terms of function value and FW gap convergence, and sparsity of the obtained solutions, compared to the standard and lazified versions (that is, using the active set vertices as a weak separation oracle (Braun et al.,, 2017)) of AFW and BPFW, which also produce notably sparse solutions. The implementation of further PM variants is left for future research. Our algorithm is implemented in Julia (Bezanson et al.,, 2017) v1.9.2 and builds on the FrankWolfe.jl package (Besançon et al.,, 2022).
@@ E Numerical experiments (extended)
E.1 Sparse logistic regression
@@
We also study and compare the effect of the pivoting meta algorithm on the lazified and non-lazified version of AFW, illustrated on another regression on the $K$-sparse polytope in Figure 4. On this instance, the pivoting technique applied to the lazified AFW algorithm only slightly improves the active set cardinality and does not change convergence. Pivoting however sharply reduces the cardinality of the active set for the non-lazified variant. This however comes at the expense of a higher per-iteration runtime and higher numerical instabilities which affect in particular the FW gap convergence. Such artifacts occur when the determinant of the matrix $M$ increases with new vertices, and the phenomenon is particularly pronounced for non-lazified variants which typically introduce more vertices. We discuss numerical robustness and stability in more detail in Subsection D.
E.2 Sparse signal recovery

### v3

@@ 8 FW variants
4 ${\mathbf{a}}^{(t)}\leftarrow\operatorname{argmax}_{{\mathbf{v}}\in{\mathcal{S}}^{(t)}}\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{v}}\rangle$ $\triangleright$ away vertex
5 ${\mathbf{v}}^{(t)}\leftarrow\operatorname{argmin}_{{\mathbf{v}}\in{\mathcal{V}}}\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{v}}\rangle$ $\triangleright$ FW vertex
@@
7    ${\mathbf{d}}^{(t)}\leftarrow{\mathbf{x}}^{(t)}-{\mathbf{a}}^{(t)}$ $\triangleright$ away step
@@
[…]
@@
12    ${\mathbf{d}}^{(t)}\leftarrow{\mathbf{v}}^{(t)}-{\mathbf{x}}^{(t)}$ $\triangleright$ FW step
@@
[…]
@@ 17 ${\mathbf{x}}^{(t+1)}\leftarrow{\mathbf{x}}^{(t)}+\eta^{(t)}{\mathbf{d}}^{(t)}$
4 ${\mathbf{a}}^{(t)}\leftarrow\operatorname{argmax}_{{\mathbf{v}}\in{\mathcal{S}}^{(t)}}\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{v}}\rangle$ $\triangleright$ away vertex
5 ${\mathbf{w}}^{(t)}\leftarrow\operatorname{argmin}_{{\mathbf{v}}\in{\mathcal{S}}^{(t)}}\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{v}}\rangle$ $\triangleright$ local FW vertex
6 ${\mathbf{v}}^{(t)}\leftarrow\operatorname{argmin}_{{\mathbf{v}}\in{\mathcal{V}}}\langle\nabla f({\mathbf{x}}^{(t)}),{\mathbf{v}}\rangle$ $\triangleright$ FW vertex
@@
8    ${\mathbf{d}}^{(t)}\leftarrow{\mathbf{w}}^{(t)}-{\mathbf{a}}^{(t)}$ $\triangleright$ local pairwise step
@@
13    ${\mathbf{d}}^{(t)}\leftarrow{\mathbf{v}}^{(t)}-{\mathbf{x}}^{(t)}$ $\triangleright$ FW step
@@
[…]
@@ 18 ${\mathbf{x}}^{(t+1)}\leftarrow{\mathbf{x}}^{(t)}+\eta^{(t)}{\mathbf{d}}^{(t)}$
9 Active set identification with the blended pairwise Frank-Wolfe algorithm
@@
[…]
@@
[…]
$\delta_{\min}=\min_{i\in I^{c}({\mathbf{x}}^{*})}\lambda_{i}({\mathbf{x}}^{*}),$ (9.2)
@@
Lemma 9.1 (BPFW step multipliers).
@@
Lemma 9.2 (Bounds on the multipliers).
@@
$\lambda_{i}({\mathbf{x}}^{(t)})$ $<h^{*}\left(L+\frac{\delta^{(t)}}{2}\right)$ $\mathrm{for\ all}\ i\in I({\mathbf{x}}^{*})$ (9.3)
$\lambda_{j}({\mathbf{x}}^{(t)})$ $>-h^{*}\left(L+\frac{\delta^{(t)}}{2}\right)+\delta_{\min}$ $\mathrm{for\ all}\ j\in I^{c}({\mathbf{x}}^{*}).$ (9.4)
@@
Bomze et al., (2020, Lemma 3.1) proves that $|\lambda_{i}({\mathbf{x}}^{*})-\lambda_{i}({\mathbf{x}}^{(t)})|<h^{*}\left(L+\frac{\delta^{(t)}}{2}\right)$ for any $i\in[n]$ and iteration $t$ where the assumptions of the Lemma hold. The two subsequent inequalities follow from the definitions of $I({\mathbf{x}}^{*})$ in (9.1) and $\delta_{\min}$ in (9.2). In particular, Equation (9.4) can be obtained as follows:
@@
$\eta^{(t)}\geq\min\left\{\alpha_{{\mathbf{a}}^{(t)}}^{(t)},c\frac{\langle\nabla f({\mathbf{x}}^{(t)}),-{\mathbf{d}}^{(t)}\rangle}{L\|{\mathbf{d}}^{(t)}\|_{2}^{2}}\right\}$ (9.5)
for a given $c>0$ is a sufficient condition for Assumption 1 to hold (with a modified Lipschitz constant $L/c$). This means in particular that the commonly-used adaptive step-size strategy introduced in Pedregosa et al., (2018) respects the weaker condition (9.5), with the factor $c$ corresponding to the parameter $\tau$ of the backtracking subroutine Pedregosa et al., (2018, Algorithm 2).
We can now state Theorem 9.3, which proves the active set identification for BPFW. The proof follows the structure of that of Bomze et al., (2020, Theorem 3.3), where the same property is shown for AFW.
Theorem 9.3 (BPFW identification).
@@
[…]
@@
Since $J^{(t)}=\emptyset$, it holds that $I^{c}({\mathbf{x}}^{*})=O^{(t)}$ and $x^{(t)}_{j}=0$ for all $j\in I^{c}({\mathbf{x}}^{*})$. Following Lemma 9.1, Property 2, we conclude that $x_{j}^{(t+1)}=0$ for all $j\in I^{c}({\mathbf{x}}^{*})$. Thus, $J^{(t+1)}=\emptyset$.
[…]
@@
We next demonstrate that the FW vertex is always selected in $I({\mathbf{x}}^{*})$ and not in $J^{(t)}$. Based on (9.3) and (9.4), a sufficient condition for the FW vertex to be selected in $I({\mathbf{x}}^{*})$ is:
@@
$\geq\frac{\delta_{\min}-h^{*}(2L+\delta^{(t)})}{2L}$ $\triangleright$ by subtracting (9.3) from (9.4)
@@
Theorem 9.3 can be used to generalize Bomze et al., (2020, Theorem 4.3) to BPFW under similar assumptions, meaning that BPFW identifies the active set in a finite number of iterations, since the proof of Bomze et al., (2020, Theorem 4.3) does not rely on the sequence generated by the algorithm. We also note that our proof does not extend to the pairwise Frank-Wolfe algorithm (Lacoste-Julien and Jaggi,, 2015), since we rely on the BPFW property that at any iteration in which a pairwise step is performed, both the away and the FW vertex are selected from the active set.
Theorem 9.4 (Active set identification property of BPFW).
@@
By affine invariance of BPFW with line-search, this is equivalent to optimizing $f(A{\mathbf{y}})$, with ${\mathbf{y}}\in\Delta_{n}$, as developed in Bomze et al., (2020, Appendix C). The proof over the simplex follows that of Bomze et al., (2020, Theorem 4.3), relying on Theorem 9.3 to ensure the cardinality of the set $J^{(t)}$ decreases at each iteration once there exists a stationary point $\mathbf{x}$$^{*}$ of the problem with ${\mathbf{x}}^{*}\in{\mathcal{X}}$, such that $\|{\mathbf{x}}^{(t)}-{\mathbf{x}}^{*}\|_{1}<h^{*}$. ∎
10 Algorithmic aspects
@@
11 Numerical experiments (extended)
We assess our algorithm on efficiency in terms of function value and FW gap convergence, and sparsity of the obtained solutions, compared to the standard and lazified versions (that is, using the active set vertices as a weak separation oracle (Braun et al.,, 2017)) of AFW and BPFW, which also produce notably sparse solutions. The implementation of further PM variants is left for future research. Our algorithm is implemented in Julia (Bezanson et al.,, 2017) v1.9.2 and builds on the FrankWolfe.jl package (Besançon et al.,, 2022). The code for the algorithm and experiments is available on the repository github.com/ZIB-IOL/pivoting_frankwolfe.
@@ 11 Numerical experiments (extended)
11.1 Sparse logistic regression
@@
We also study and compare the effect of the pivoting meta algorithm on the lazified and non-lazified version of AFW, illustrated on another regression on the $K$-sparse polytope in Figure 4. On this instance, the pivoting technique applied to the lazified AFW algorithm only slightly improves the active set cardinality and does not change convergence. Pivoting however sharply reduces the cardinality of the active set for the non-lazified variant. This however comes at the expense of a higher per-iteration runtime and higher numerical instabilities which affect in particular the FW gap convergence. Such artifacts occur when the determinant of the matrix $M$ increases with new vertices, and the phenomenon is particularly pronounced for non-lazified variants which typically introduce more vertices. We discuss numerical robustness and stability in more detail in Subsection 10.
11.2 Sparse signal recovery